OpenStudy (anonymous):

f(x)=x^(3)/(x^(2)-4) what interval contains a local maximum for the function above.

OpenStudy (anonymous):

\[f(x)=\frac{x^3}{(x+2)(x-2)}\] ?

OpenStudy (anonymous):

\[f(x)=x^3/(x^2-4)\]

OpenStudy (anonymous):

same thing calc class or algebra?

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

ok you need to graph then you have to vertical asymptotes at \(x=-2\) and \(x=2\)

OpenStudy (anonymous):

in the middle it goes from \(\infty\) to \(-\infty\) so no max on \((-2,2)\)

OpenStudy (anonymous):

to the left of \(-2\) it goes up and then down, so that is the interval that will have a local max to the right of \(2 \) it goes down and then up, that is the interval that will have a local min

OpenStudy (anonymous):

A. (0,2) B. (-(infinity), -2) C.(-2,0) D.(2, (infinty) These are my answer choices

OpenStudy (anonymous):

did you understand what i wrote? the answer is there

OpenStudy (anonymous):

no i do not understand it.

OpenStudy (anonymous):

(-2,0)?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

no the interval will be \((2,\infty)\) on which there will be a minimum

OpenStudy (anonymous):

here is a picture, you can see why this is true http://www.wolframalpha.com/input/?i= \frac{x^3}{%28x%2B2%29%28x-2%29}